The leader of the jeepney drivers’ association claims that the average daily take-home pay of all jeepney drivers in Pasay city is Php 400.00. A random sample of 100 jeepney drivers in Pasay city was interviewed and the average daily take-home pay of these drivers is found to be Php 425. Use a 0.05 level of significance and assume the standard deviation is Php 92.

 Ծ = 92          Ẋ = 425           ﬠ = 400           n = 100         level of significance = 0.05

 

1. State the hypothesis

  Ho: u = 400

  Ha: u ‡ 400

 

2. Type of test: Two-tailed test

 

 

3. Critical Value

 

 

4. Test statistics

 

 

5. Reject or accept the null hypothesis


1
Expert's answer
2022-05-30T03:34:04-0400

1. State the hypothesis

H0"\mu=400"

Ha"\mu\ne400"


2. Type of test: Two-tailed test

 

 

3. Critical Value

 "\alpha=0.05, \alpha/2=0.025, z_{0.025}=1.96" (using t-table)

 So, the rejection region for this two-tailed test is "R=\left\{z:∣z∣>1.96\right\}."


4. Test statistics

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{425-400}{92/\sqrt{100}}=2.72"


5. Reject or accept the null hypothesis

Since 1.96<2.72, "z_{0.025}<z" . So z is at rejection region.

Thus, "H_0" should be rejected. The average daily take home pay of all jeepney drivers in Pasay City is different from P400.00